ID |
Date |
Author |
Type |
Category |
Subject |
16547
|
Thu Jan 6 13:54:28 2022 |
Koji | Update | CDS | Yearly DAQD fix 2022! |
Just restarting all the c1sus2 models fixed the issue. (Attachment 1)
SUS2 ADC1 CH21 is saturated. I'm not yet sure if this is the electronics issue or the ADC issue.
SUS2 ADC1 CH10 also has large offset. This should also be investiagted. |
Attachment 1: Screenshot_2022-01-06_13-57-40.png
|
|
16546
|
Thu Jan 6 12:52:49 2022 |
Anchal | Update | CDS | Yearly DAQD fix 2022! |
Just as predicted, all realtime models reported "0x4000" error. Read the parent post for more details. I fixed this by following the instructions. I add folowing lines to the file /opt/rtcds/rtscore/release/src/include/drv/spectracomGPS.c in fb1:
/* 2020 had 366 days and no leap second */
pHardware->gpsOffset += 31622400;
/* 2021 had no leap seconds or leap days, so adjust for that */
pHardware->gpsOffset += 31536000;
Then is made the package and reloaded it after stoping the daqd services. This brought back all the fast models except C1SUS2 models which are in red due to some other reason that I'll investigate further.
|
16545
|
Thu Jan 6 11:54:20 2022 |
Anchal | Summary | BHD | Part IX of BHR upgrade - Placed AS1 and AS4 |
[Paco (Vacuum Work), Anchal]
Today we opened the ITMY Chamber and installed suspended AS1 and AS4 in their planned positions. In doing so, we removed the razor or plate mounted on a pico motor at the south end of the table (see 40m/16450). We needed to make way for AS4 to be installed.
Photos: https://photos.app.goo.gl/YP2ZZhQ3jip3Uhp5A
We need more dog clamps for installing the suspensions, we have used temporary clamps for now. However, knows where new C&B clamps are, please let us know. |
16544
|
Wed Jan 5 19:18:06 2022 |
Yehonathan | Update | BHD | SOS assembly -- AS4 |
{Paco, Yehonathan, Anchal}
Today we suspended AS4 (E2000226-B). Anchal mounted Lambda Optic mirror with an RoC closest to AS4 in a thin optic mount. He noted that this optic as well as AS1 don't have a wedge angle. The specs claim that the wedge angle is 2 degrees what should have been clearly seen by inspecting the optic with a naked eye. All the ghost beam deflections probably come from the curvature of the mirror.
We did all the height and roll balancing using a camera (Attachment 1,2). We balanced that pitch of the adapter using a QPD not before we realigned the OpLev setup.
We measured the motion spectra (attachment 3). Major peaks are found at 755 mHz, 964 mHz, and 1.062Hz. I locked the counterweights setscrew and observed that the pitch balance doesn't change. I locked the EQ stops such that the alignment of the mirror remained the same by monitoring the QPD signals. I clamped the suspensions wires to the suspension block.
The only thing remaining is inserting the OSEMs.
|
Attachment 1: AS4_roll_balance.png
|
|
Attachment 2: AS_4_magnet_height.png
|
|
Attachment 3: FreeSwingingSpectra.pdf
|
|
16543
|
Wed Jan 5 17:46:04 2022 |
Anchal | Update | BHD | Tested 2" PR2 candidates transmission |
I tested 2 more optics today, the old PR2 that we took out and another optic I found in QIL. Both these optics are also not good for our purpose.
|
Polarization |
Incident Power [mW] |
Transmitted Power [mW] |
Transmission [ppm] |
Existing PR2 |
p-pol |
910 |
0.004 |
4.4 |
V2-1698 & V2-1700 |
p-pol |
910 |
595 |
653846 |
I'll find thw Y1S optic and test that too. We should start looking for alternate solutions as well.
|
16542
|
Tue Jan 4 18:27:23 2022 |
Paco | Update | BHD | SOS assembly -- PR3 |
[yehonathan, paco, anchal]
We continue suspending PR3 today. Yehonathan and Paco suspended the thick optic in its adapter. After fixing some nominal height and undoing any residual roll angle (see Attachments 1,2 for pictures), we noticed a problem with the pitch angle, so we insert the counterweights all the way in. Nevertheless, we soon found out that we needed to shift one of the two counterweights to the back of the adapter side (so one on each side) in order to tare the pitch angle. This is a newly experienced maneuver that may apply for further thick optics.
After taring the pitch angle roughly, we noted another issue. The wedge (~ 1 deg) on the optic made it such that the protruding socket heads on the thick side bumped against the lower clamp (not the earthquake stop tip itself). Attachments #4,5 show the before/after situation which was solved provisionally by replacing the socket head screws with lower profile (flat) head screws in situ. Again, this operation was highly delicate and specific to wedged thick optics, so for future SOS we should keep it in mind.
Another issue that we had with the new thick optic adapters is that for some reason there is a recession in the upper backside of the adapter (attachment coming soon). This makes the upper back EQ stop too short to touch the adapter. We replaced it with a longer screw. When inserted it doesn't really hit the back of the adapter. Rather, it touches the corner of the recession, stoping the optic with friction.
While all this was happening, Anchal started mounting AS4 on its adapter. After one of the magnets broke off, he switched to another one and succeeded. This is the next target for suspension. We still need to check the orientation of the wedge. Furthermore, we started a gluing session in the afternoon to prepare as much as possible for further SOS during the week. 3 side magnets were glued to side blocks. 3 magnets were glued to 3 adapters that were missing 1 magnet each.
In the afternoon, Yehonathan and Paco set up the QPD and did all the usual balancing, and then Anchal took the data of which the result is shown in Attachment #3. The major peaks are located at 723mHz, 953mHz, and 1.05Hz. Very similar to the case of the thin optic adapters.
Anchal progressed with OSEM installation, and engraving and yehonathan glued the counterweight setscrew in place. After securing the EQ stops, and wrapping the wires in foil, we declare PR3 is ready to be installed. |
Attachment 1: PR3_roll_balance.png
|
|
Attachment 2: PR3_magnet_height.png
|
|
Attachment 3: FreeSwingingSpectra.pdf
|
|
Attachment 4: PXL_20220104_231742123.jpg
|
|
Attachment 5: PXL_20220104_232809203.jpg
|
|
16541
|
Tue Jan 4 18:26:59 2022 |
Anchal | Update | BHD | Tested 2" PR2 candidates transmission |
I used the rejected light from the PBS after the motorized half-wave plate between PMC and IMC injection path (used for input power control to IMC) to measure the transmission of PR2 candidates. These candidates were picked from QIL (QIL/2696). Unfortunately, I don't think either of these mirrors can be used for PR2.
|
Polarization |
Incident Power [mW] |
Transmitted Power [mW] |
Transmission [ppm] |
V2-2239 & V2-2242 |
s-pol |
940 |
0.015 |
16.0 |
V2-2239 & V2-2242 |
p-pol |
935 |
0.015 |
16.0 |
V6-704 & V6-705 |
p-pol |
925 |
21 |
22703 |
If I remember correctly, we are looking for a 2" flat mirror with a transmission of the order of 1000 ppm. The current PR2 is supposed to have less than 100 ppm transmission which would not leave enough light for LO path.
I've kept the transmission testing setup intact on the PSL table, I'll test existing PR2 and another optic (which is 0.5" thick unfortunately) tomorrow. |
16540
|
Mon Jan 3 16:46:41 2022 |
Paco | Update | BHD | 1Y1 rack work for SR2, PR2, PR3 |
[Paco, Anchal]
Continued working on 1Y1 rack. Populated the 6 coil drivers, made all connections between sat amp, AA chassis, DAC, and ADC adapters for SR2, PR2, and PR3 suspensions. Powered all boxes and labeled them and cables where needed. Near the end, we had to increase the current limit on the positive rail sorensen (+18 V) from ~ 7 to > 8.0 Amps to feed all the instruments. We also increased the negative (-18 V) current limit proportionally.
We think we are ready for all the new SOS on this side electronics-wise.
Photos: https://photos.app.goo.gl/GviuqLQviSPo1M3G6 |
16539
|
Mon Jan 3 12:05:08 2022 |
Paco | Update | BHD | 1Y0 rack work for LO2 AS1 AS4 |
[Paco, Anchal]
Continue working on 1Y0. Added coil drivers for LO2, AS1, AS4. Anchal made additional labels for cables and boxes. We lined up all cables, connected the different units and powered them without major events. |
16538
|
Sun Jan 2 20:46:46 2022 |
Koji | Update | SUS | End SUS Electronics building |
19:00~ Start working on the electronics bench
The following units were tested and ready to be installed. These are the last SUS electronics units and we are now ready to upgrade the end SUS electronics too.
40m End ADC Adapter Unit D2100016 / 2 Units (S2200001 S2200002)
40m End DAC Adapter Unit D2100647/ 2 Units (S2200003 S2200004)
These are placed on Tega's desk together with the vertex DAC adapters
0:30 End work |
Attachment 1: PXL_20220103_081133119.jpg
|
|
16537
|
Wed Dec 29 20:09:40 2021 |
rana | Summary | CDS | c1su2 model updated with SUS damping blocks for 7 SOSs |
We want to maintain the 16 kHz sample rate for the COIL DAQ channels, but nothing wrong with reducing the others.
I would suggest setting the DQ sample rates to 256 Hz for the SUS DAMP channels and 1024 Hz for the OPLEV channels (for noise diagnostics).
Maybe you can put these numbers into a new library part and we can have the best of all worlds?
Quote: |
Should we change the library model part for sus_single_control.mdl
We notice that all our suspension models need to go through this weird python script modifying auto-generated .ini files to reduce the data rate. Ideally, there is a simpler solution to this by simply adding the datarate 2048 in the '#DAQ Channels' block in the model library part /cvs/cds/rtcds/userapps/trunk/sus/c1/models/lib/sus_single_control.mdl which is the root model in all the suspensions. With this change, the .ini files will automatically be written with correct datarate and there will be no need for using the activateDQ script. But we couldn't find why this simple solution was not implemented in the past, so we want to know if there is more stuff going on here then we know. Changing the library model would obviously change every suspension model and we don't want a broken CDS system on our head at the begining of holidays, so we'll leave this delicate task for the near future.
|
|
16536
|
Fri Dec 24 16:49:41 2021 |
Koji | Update | General | Is megatron down? (Re: chiara local backup) |
It turned out that the UPS installed on Nov 22 failed (cf https://nodus.ligo.caltech.edu:8081/40m/16479 ). As a fact, it was alive just for 2 weeks!
The APC UPS unit indicated F06. According to the manual (https://www.apc.com/shop/us/en/products/APC-Power-Saving-Back-UPS-Pro-1000VA/P-BR1000G), F06 means "Relay Welding" and can not be fixed by a user. Resetting the UPS eliminated the error, but I didn't want to have the same issue while no one is in the lab, I moved the megatron power source from the UPS to the power strip on 1Y7. So, megatron is currently vulnerable to a power glitch.
After the power cords were restored, megatron eventually recovered ssh terminals. I manually ran autoburt.cron at 16:50 so that the latest snapshot is taken. |
Attachment 1: PXL_20211224_235652821.jpg
|
|
16535
|
Thu Dec 23 16:38:21 2021 |
Koji | Update | General | Is megatron down? (Re: chiara local backup) |
The local backup seems working fine again. But I found that megatron is down and this is a real issue. This should be fixed at the earliest chance.
It seems that the local backup has been successfully taken this morning.
controls@nodus|backup> tail /opt/rtcds/caltech/c1/scripts/backup/localbackup.log
2021-12-19 07:00:01,146 INFO Updating backup image of /cvs/cds
2021-12-19 07:00:01,146 ERROR External drive not mounted!!!
2021-12-20 07:00:01,255 INFO Updating backup image of /cvs/cds
2021-12-20 07:00:01,255 ERROR External drive not mounted!!!
2021-12-21 07:00:01,361 INFO Updating backup image of /cvs/cds
2021-12-21 07:00:01,361 ERROR External drive not mounted!!!
2021-12-22 07:00:01,469 INFO Updating backup image of /cvs/cds
2021-12-22 07:00:01,470 ERROR External drive not mounted!!!
2021-12-23 07:00:01,594 INFO Updating backup image of /cvs/cds
2021-12-23 07:19:55,560 INFO Backup rsync job ran successfully, transferred 338425 files.
However, I noticed that the autoburt has been stalled since Dec 6 (I used to check how the backup is up-to-date using the autoburt snapshots)
Dec>pwd
/opt/rtcds/caltech/c1/burt/autoburt/snapshots/2021/Dec
Dec>ls -l
total 24
drwxr-xr-x 26 controls controls 4096 Dec 1 23:07 1
drwxr-xr-x 26 controls controls 4096 Dec 2 23:07 2
drwxr-xr-x 26 controls controls 4096 Dec 3 23:07 3
drwxr-xr-x 26 controls controls 4096 Dec 4 23:07 4
drwxr-xr-x 26 controls controls 4096 Dec 5 23:07 5
drwxr-xr-x 19 controls controls 4096 Dec 6 16:07 6
There are a bunch of errors in the log file as follows, but maybe this is not an issue
controls@nodus|burt> pwd
/opt/rtcds/caltech/c1/burt
controls@nodus|burt> tail burtcron.log
!!! ERROR !!! Target c1supepics Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1tstepics Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1x10epics Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1aux Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1dcuepics Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1iscaux Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1iscepics Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1losepics Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1psl Snapshot file inconsistent with Request file
!!! ERROR !!! Target c1susaux Snapshot file inconsistent with Request file
The real issue seems that megatron is down. It has a lot of house keeping jobs on corn including the N2 pressure alert.
https://wiki-40m.ligo.caltech.edu/Computers_and_Scripts/CRON
This needs to be fixed at the earliest chance. |
16534
|
Wed Dec 22 18:16:23 2021 |
Koji | Update | SUS | Remaining task for 2021 |
The in-vacuum installation team has reported that the side OSEMs of ITMX and LO1 are going to be interfering if place LO1 at the planned location.
I confirmed that ITMX has the side magnet on the other side (Attachment 1 ITMX photo taken on 2016/7/21). So we can do this swap.
The ITMX side OSEM is sticking out most. By doing this operation, we will recover most of the space between the ITMX and LO1. (Attachment 2) |
Attachment 1: ITMX_2016_07_21.jpg
|
|
Attachment 2: Screen_Shot_2021-12-22_at_18.03.42.png
|
|
16533
|
Wed Dec 22 17:40:22 2021 |
Anchal | Summary | CDS | c1su2 model updated with SUS damping blocks for 7 SOSs |
[Anchal, Koji]
I've updated the c1su2 model today with model suspension blocks for the 7 new SOSs (LO1, LO2, AS1, AS4, SR2, PR2 and PR3). The model is running properly now but we had some difficulty in getting it to run.
Initially, we were getting 0x2000 error on the c1su2 model CDS screen. The issue probably was high data transmission required for all the 7 SOSs in this model. Koji dug up a script /opt/rtcds/caltech/c1/userapps/trunk/cds/c1/scripts/activateDQ.py that has been used historically for updating the data rate on some of theDQ channels in the suspension block. However, this script was not working properly for Koji, so he create a new script at /opt/rtcds/caltech/c1/chans/daq/activateSUS2DQ.py.
[Ed by KA: I could not make this modified script run so that I replaces the input file (i.e. C1SU2.ini). So the output file is named C1SU2.ini.NEW and need to manually replace the original file.]
With this, Koji was able to reduce acquisition rate of SUSPOS_IN1_DQ, SUSPIT_IN1_DQ, SUSYAW_IN1_DQ, SUSSIDE_IN1_DQ, SENSOR_UL, SENSOR_UR, SENSOR_LL,SENSOR_LR, SENSOR_SIDE, OPLEV_PERROR, OPLEV_YERROR, and OPLEV_SUM to 2048 Sa/s. The script modifies the /opt/rtcds/caltech/c1/chans/daq/C1SU2.ini file which would get re-written if c1su2 model is remade and reinstalled. After this modification, the 0x2000 error stopped appearing and the model is running fine.
Should we change the library model part for sus_single_control.mdl
We notice that all our suspension models need to go through this weird python script modifying auto-generated .ini files to reduce the data rate. Ideally, there is a simpler solution to this by simply adding the datarate 2048 in the '#DAQ Channels' block in the model library part /cvs/cds/rtcds/userapps/trunk/sus/c1/models/lib/sus_single_control.mdl which is the root model in all the suspensions. With this change, the .ini files will automatically be written with correct datarate and there will be no need for using the activateDQ script. But we couldn't find why this simple solution was not implemented in the past, so we want to know if there is more stuff going on here then we know. Changing the library model would obviously change every suspension model and we don't want a broken CDS system on our head at the begining of holidays, so we'll leave this delicate task for the near future. |
16532
|
Wed Dec 22 14:57:05 2021 |
Koji | Update | General | chiara local backup |
chiara local backup of /cvs/cds has not been running since the move of chiara in Nov 19. The remote backup has not been taken since 2017.
The lack of the local backup was because of the misconfiguration of /etc/fstab.
It was fixed and now the backup disk was mounted. We'll see the backup script running tomorrow morning.
The backup disk is smaller than the main disk. So sooner or later, we will face the backup problem again.
localbackup script was crying because there was no backup disk.
backup>pwd
/opt/rtcds/caltech/c1/scripts/backup
backup>tail localbackup.log
2021-12-18 07:00:02,002 INFO Updating backup image of /cvs/cds
2021-12-18 07:00:02,002 ERROR External drive not mounted!!!
2021-12-19 07:00:01,146 INFO Updating backup image of /cvs/cds
2021-12-19 07:00:01,146 ERROR External drive not mounted!!!
2021-12-20 07:00:01,255 INFO Updating backup image of /cvs/cds
2021-12-20 07:00:01,255 ERROR External drive not mounted!!!
2021-12-21 07:00:01,361 INFO Updating backup image of /cvs/cds
2021-12-21 07:00:01,361 ERROR External drive not mounted!!!
2021-12-22 07:00:01,469 INFO Updating backup image of /cvs/cds
2021-12-22 07:00:01,470 ERROR External drive not mounted!!!
fstab had no entry for the backup disk.
backup>cat /etc/fstab
# /etc/fstab: static file system information.
#
# Use 'blkid -o value -s UUID' to print the universally unique identifier
# for a device; this may be used with UUID= as a more robust way to name
# devices that works even if disks are added and removed. See fstab(5).
#
# <file system> <mount point> <type> <options> <dump> <pass>
proc /proc proc nodev,noexec,nosuid 0 0
# / was on /dev/sda1 during installation
UUID=972db769-4020-4b74-b943-9b868c26043a / ext4 errors=remount-ro 0 1
# swap was on /dev/sda5 during installation
UUID=a3f5d977-72d7-47c9-a059-38633d16413e none swap sw 0 0
# OLD BACKUP DISK
#UUID="90a5c98a-22fb-4685-9c17-77ed07a5e000" /media/40mBackup ext4 defaults,relatime,commit=60 0 0
# CURRENT BACKUP DISK as of 2021/09/02
#UUID="1843f813-872b-44ff-9a4e-38b77976e8dc" /media/40mBackup ext4 defaults,relatime,commit=60 0 0
#fb:/frames /frames nfs ro,bg
# CURRENT MAIN DISK as of 2021/09/02
# UUID=92dc7073-bf4d-4c58-8052-63129ff5755b /home/cds ext4 defaults,relatime,commit=60 0 0
UUID="1843f813-872b-44ff-9a4e-38b77976e8dc" /home/cds ext4 defaults,relatime,commit=60 0 0
Checked the dev name of the disks and the UUIDs
backup>sudo lsblk
[sudo] password for controls:
NAME MAJ:MIN RM SIZE RO TYPE MOUNTPOINT
sda 8:0 0 465.8G 0 disk
├─sda1 8:1 0 446.9G 0 part /
├─sda2 8:2 0 1K 0 part
└─sda5 8:5 0 18.9G 0 part [SWAP]
sdb 8:16 0 5.5T 0 disk
└─sdb1 8:17 0 5.5T 0 part /home/cds
sdc 8:32 0 3.7T 0 disk
└─sdc1 8:33 0 3.7T 0 part
sr0 11:0 1 1024M 0 rom
backup> sudo blkid
/dev/sda1: UUID="972db769-4020-4b74-b943-9b868c26043a" TYPE="ext4"
/dev/sda5: UUID="a3f5d977-72d7-47c9-a059-38633d16413e" TYPE="swap"
/dev/sdb1: UUID="1843f813-872b-44ff-9a4e-38b77976e8dc" TYPE="ext4"
/dev/sdc1: UUID="92dc7073-bf4d-4c58-8052-63129ff5755b" TYPE="ext4"
Added the fstab entry for the backup disk
media>cat /etc/fstab
# /etc/fstab: static file system information.
#
# Use 'blkid -o value -s UUID' to print the universally unique identifier
# for a device; this may be used with UUID= as a more robust way to name
# devices that works even if disks are added and removed. See fstab(5).
#
# <file system> <mount point> <type> <options> <dump> <pass>
proc /proc proc nodev,noexec,nosuid 0 0
# / was on /dev/sda1 during installation
UUID=972db769-4020-4b74-b943-9b868c26043a / ext4 errors=remount-ro 0 1
# swap was on /dev/sda5 during installation
UUID=a3f5d977-72d7-47c9-a059-38633d16413e none swap sw 0 0
# OLD BACKUP DISK
#UUID="90a5c98a-22fb-4685-9c17-77ed07a5e000" /media/40mBackup ext4 defaults,relatime,commit=60 0 0
# OLD BACKUP DISK as of 2021/09/02
#UUID="1843f813-872b-44ff-9a4e-38b77976e8dc" /media/40mBackup ext4 defaults,relatime,commit=60 0 0
# Current backup disk as of 2021/12/22
UUID="92dc7073-bf4d-4c58-8052-63129ff5755b" /media/40mBackup ext4 defaults,relatime,commit=60 0 0
#fb:/frames /frames nfs ro,bg
# CURRENT MAIN DISK as of 2021/09/02
# UUID=92dc7073-bf4d-4c58-8052-63129ff5755b /home/cds ext4 defaults,relatime,commit=60 0 0
UUID="1843f813-872b-44ff-9a4e-38b77976e8dc" /home/cds ext4 defaults,relatime,commit=60 0 0
|
16531
|
Tue Dec 21 18:04:46 2021 |
Yehonathan | Update | BHD | SOS assembly |
I locked the EQ stops while retaining the XY alignment on the QPD and installed 5 green OSEMs. AS1 is ready for transfer into the vacuum chamber. |
16530
|
Tue Dec 21 16:52:39 2021 |
Anchal | Summary | Electronics | In-air Sat Amp to Vacuum Flange cables laid for 7 new SOS |
[Anchal, Yehonathan, Chub]
We today laid down 14 70 ft long DB25 cables from 1Y1 (6), 1Y0 (8) to ITMY Chamber (4), BS Chamber (6) and ITMX Chamber (4). The cables have been connected to respective satellite amplifier on the racks and the other ends are connected to the vacuum flange feedthru on ITMX for LO1 and PR2, while the others have been kept near the planned flange postions. LO1 is now ready to be connected to CDS by connecting the in-vacuum cable inside ITMX chamber to the OSEMs. |
16529
|
Tue Dec 21 16:35:39 2021 |
Koji | Update | VAC | ITMX NW feedthru (LO1-1) connector pin bent |
I've received a report that a pin of an ITMX NW feedthru connector was bent. (Attachment 1)
The connector is #1 (upper left) and planned to be used for LO1-1.
This is Pin25 and used for the PD K of OSEM #1. This means that Coil Driver #1 (3 OSEMs) uses this pin, but Coil Driver #2 (2 OSEMs) does not.
Anyways, I tried to fix it by bending it back. WIth some tools, it was straightened enough for plugging the cable connector. (Attachment 2)
It seemed that the pins were exceptionally soft compared to the ones used for usual DSUBs, probably because of the vacuum compatibility.
So it's better to approach the pins in parallel to the surface and not apply mating pressure until you are sure that all the 25pins are inserted in the counterpart holes. |
Attachment 1: PXL_20211222_002019620.jpg
|
|
Attachment 2: PXL_20211222_003014068.jpg
|
|
16528
|
Mon Dec 20 17:26:13 2021 |
Yehonathan | Update | BHD | SOS assembly |
{Yehonathan, Anchal}
I released the AS1 wires from the winches, removed the adapter from the SOS tower, and removed the Lambda optic from the adapter. Attachment 1 shows the pencil markings on the optic before cleaning. I cleaned the pencil marking from the side of the optic with acetone using swabs until there were no pencil residues on the swab (attachment 2 shows the swab I used next to an unused swab). I was not able to remove the markings completely though (attachment 3).
I remounted the optic with the arrow rotated by 90 degrees counterclockwise.
We hang the adapter on the winches and adjust the height of the magnet and the adapter roll using the winches. We monitor the height of the adapter using a live stream from the Cannon camera. The camera's tilt was adjusted using straight features on the SOS tower. When we ran out of winch travel we adjust the height using the lower EQ stops and pull tight the wires. Attachment 4 shows the alignment of the side magnet with respect to the SOS tower and a side OSEM.
We checked the ghost beam trajectory and it looks much better (attachment 5)
We started realigning the OpLev. We realize that the height of the beam should be 5+14/32" = 5.437 by measuring the height of the screw holding the side OSEM from the table. The real height from the schematics is 5.425 We make the beam parallel with the table first using an iris and then the QPD.
Today, I balanced the counterweight. First using an iris, then by placing a QPD close to the SOS measuring the reflection from AS1. I locked the counterweight's set screw and the QPD Y readout looks good. Attachment 6 shows the QPD y readout near the beat node between pitch and pos. The node comes very close to zero which indicates that the pitch is balanced.
I measured the free-swinging motion using the QPD x and y axes. Attachment 7 shows the spectra of that motion. The major peaks are at 755mHz, 953mHz, and 1.05Hz.
|
Attachment 1: IMG_6312.JPG
|
|
Attachment 2: IMG_6315.JPG
|
|
Attachment 3: IMG_6314.JPG
|
|
Attachment 4: AS1adapterbalancing.png
|
|
Attachment 5: as1ghostbeam2.png
|
|
Attachment 6: beat.png
|
|
Attachment 7: FreeSwingingSpectra.pdf
|
|
16527
|
Mon Dec 20 14:10:56 2021 |
Anchal | Update | BHD | All coil drivers ready to be used, modified and tested |
Koji found some 68nF caps from Downs and I finished modifying the last remaining coil driver box and tested it.
SERIAL # |
TEST result |
S2100633 |
PASS |
With this, all coil drivers have been modified and tested and are ready to be used. This DCC tree has links to all the coil driver pages which have documentation of modifications and test data. |
16526
|
Mon Dec 20 13:52:01 2021 |
Koji | Update | BHD | SOS assembly |
LO1: No need to remove the pencil mark for the damping test. Until we see serious contamination on the LO1 optic, we don't need to take the optic off from the mount and clean it. If there is a chance of rehanging (because of a broken wire/etc), we do wipe the pencil mark.
Other optics: wipe the pencil mark as much as possible. |
16525
|
Sun Dec 19 07:52:51 2021 |
Anchal | Update | SUS | Remaining task for 2021 |
The I/O chassis reassigns the ADC and DAC indices on every power cycle. When we moved it, it must have changed it from the order we had. We were aware of this fact and decided to reconnect the I/O chassis to AA/AI to relect the correct order. We forgot to do that but this is not an error, it is expected behavior and can be solved easily.
Quote: |
I had the fear that any mistake in the electronics chain could have been the show stopper.
So I quickly checked the signal assignments for the ADC and DAC chains.
I had initial confusion (see below), but it was confirmed that the electronics chains (at least for LO1) are correct.
Note: One 70ft cable is left around the 1Y0 rack
There are a few points to be fixed:
- It looks like the ADC/DAC card # assignment has been messed up.
CDS ADC0 -> Cable label ADC1 -> AA A1 -> ...
CDS ADC1 -> Cable label ADC0 -> AA A0 -> ...
CDS DAC0 -> Cable label DAC2 -> AI D2 -> ...
CDS DAC1 -> Cable label DAC0 -> AI D0 -> ...
CDS DAC2 -> Cable label DAC1 -> AI D1 -> ...
(What is going on here... please confirm and correct as they become straight forward)
Once this puzzle was solved I could confirm reasonable connections from the end of the 70 cables to the ADC/DAC.
- We also want to change the ADC card assignment. The face OSEM readings must be assigned to ADC1 and the side OSEM readings to ADC0.
My system wiring diagram needs to be fixed accordingly too.
This is because the last channel of the first ADC (ADC0) is not available for us and is used for DuoTone.
|
|
16524
|
Sat Dec 18 00:56:14 2021 |
Koji | Update | BHD | SOS assembly |
Sad... We just need to check the wedge direction everytime, unfortunately.
Pencil: can you try to gently wipe it off with solvent & a swab? (IPA / Acetone)
If it does not come off in the end, it's all right to leave. Do we want to scribe the arrow mark? You need a diamond pen. |
16523
|
Fri Dec 17 22:16:07 2021 |
Yehonathan | Update | BHD | SOS assembly |
I specifically checked the specification before mounting the mirror. It says clearly "Arrow at the thinnest location pointing towards Side 1". I guess they just ignored it.
As for LO1, I mounted it without noticing the location of the arrow. Later, I checked and the ghost beam was horizontal so I left it as it is. Yeah, I guess I will remount the mirror. Also, what do we do with the pencil markings? It's not vacuum-compatible.
Quote: |
We @40m do the convention of the arrow at the thinnest side & pointing the HR side, but nobody says Lambda does the same.
We can just remount the mirror without breaking the wires and adjust the pitching if you do it carefully.
Does this mean that the LO1 also likely to have the wedge pointing up? Or did you rotate the mirror to have the wedge reflection to be as horizontal as possible?
|
|
16522
|
Fri Dec 17 19:19:42 2021 |
Koji | Update | SUS | Remaining task for 2021 |
I had the fear that any mistake in the electronics chain could have been the show stopper.
So I quickly checked the signal assignments for the ADC and DAC chains.
I had initial confusion (see below), but it was confirmed that the electronics chains (at least for LO1) are correct.
Note: One 70ft cable is left around the 1Y0 rack
There are a few points to be fixed:
- It looks like the ADC/DAC card # assignment has been messed up.
CDS ADC0 -> Cable label ADC1 -> AA A1 -> ...
CDS ADC1 -> Cable label ADC0 -> AA A0 -> ...
CDS DAC0 -> Cable label DAC2 -> AI D2 -> ...
CDS DAC1 -> Cable label DAC0 -> AI D0 -> ...
CDS DAC2 -> Cable label DAC1 -> AI D1 -> ...
(What is going on here... please confirm and correct as they become straight forward)
Once this puzzle was solved I could confirm reasonable connections from the end of the 70 cables to the ADC/DAC.
- We also want to change the ADC card assignment. The face OSEM readings must be assigned to ADC1 and the side OSEM readings to ADC0.
My system wiring diagram needs to be fixed accordingly too.
This is because the last channel of the first ADC (ADC0) is not available for us and is used for DuoTone. |
Attachment 1: PXL_20211218_030145356.MP.jpg
|
|
16521
|
Fri Dec 17 19:16:45 2021 |
Koji | Update | BHD | SOS assembly |
We @40m do the convention of the arrow at the thinnest side & pointing the HR side, but nobody says Lambda does the same.
We can just remount the mirror without breaking the wires and adjust the pitching if you do it carefully.
Does this mean that the LO1 also likely to have the wedge pointing up? Or did you rotate the mirror to have the wedge reflection to be as horizontal as possible? |
16520
|
Fri Dec 17 17:50:17 2021 |
Yehonathan | Update | BHD | SOS assembly |
I threaded a new wire through a different side block with a magnet and clamped it under a microscope. It was hard, but eventually, I was able to do it by holding the wire on both sides of the side block with weights.
The dented wire was discarded and the side block that was mounted on the AS1 adapter was put aside. I mounted the side block with the new wire on the AS1 adapter.
Anchal and I hanged the AS1 adapter and clamped the wires on the winches of an SOS tower. I balanced the roll and adjusted the height of the magnet with respect to a side OSEM using a camera (attachments 1 & 2).
I shoot the Hene laser on the optic and look at the reflection. I align the laser beam to be as close as possible to the center of the mirror. The OpLev needs to be realigned.
To my surprise, the ghost beam shoots up above the reflected beam! See attachment 3. I check to see that the arrow which marks the thinnest side of the mirror is horizontal (attachment 4). WTF?!
Also, now I realize that the marking on the Lambda optics are pencil markings 😵😵😵. |
Attachment 1: AS1rollbalance.png
|
|
Attachment 2: AS1Magnet_height.png
|
|
Attachment 3: ghostbeam.png
|
|
Attachment 4: lambdaopticarrow.jpg
|
|
16519
|
Fri Dec 17 12:32:35 2021 |
Koji | Update | SUS | Remaining task for 2021 |
Anything else? Feel free to edit this entry.
- SUS: AS1 hanging
- SUS: PR3/SR2/LO2 3/4" thick optic hanging
v Electronics chain check (From DAC to the end of the in-air cable / From the end of the in-air cable to the ADC)
[ELOG 16522]
- Long cable installation (4x 70ft)
- In-air cable connection to the flange
- In-vac wiring (connecting LO1 OSEMs)
- LO1 OSEM insertion/alignment
- LO1 Damping test
|
16518
|
Thu Dec 16 18:16:36 2021 |
Yehonathan | Update | BHD | SOS assembly |
Today I glued magnets onto the new 3/4" mirror adapters. I also took the opportunity to make some more side magnets assemblies.
Yesterday I mounted PR3/SR2 3/4" thick mirror onto one of the new adapter. There seem to be no issues for now.
I started the process of suspending AS1 (E2000226-A). The Lambda Optic mirror with the closest specs has Rc = 2 m. I attached side blocks with clamped wires onto adapter number 7 - side block with a magnet on the right.
I then took one of the Lambda Optic mirrors and tried mounting it in the adapter. It was quite difficult to get it right. Unfortunately, I chipped the edge of the substrate (attachment 1) 🤦🏻♂️. I put the mirror back in the box and decided to use the spare mirror. I successfully mounted it into the adapter but when I put the clamping screws one of them fell on the mirror 🤦🏻♂️🤦🏻♂️. There is no visible damage though. I took some pictures (attachment 2-4).
I and Anchal then started suspending the mirror but then we found that one of the wires is dented in the middle 🤦🏻♂️🤦🏻♂️🤦🏻♂️. I'm burned out for today.
Late update: one nice thing that I found yesterday is that the glue is viscous enough to hold the dumbells without a metal sheet from above holding the magnets. This greatly simplifies the gluing process.
|
Attachment 1: chippedmirror.png
|
|
Attachment 2: IMG_6311.JPG
|
|
Attachment 3: IMG_6310.JPG
|
|
Attachment 4: IMG_6309.JPG
|
|
16517
|
Thu Dec 16 17:57:17 2021 |
Anchal | Update | BHD | Finished Coil driver (odd serial number) units tests |
S2100619 was fixed by Koji and it passed the test after that.
|
16516
|
Thu Dec 16 17:41:12 2021 |
Koji | Update | BHD | Coil driver test failed for S2100619-v1 |
Good catch. It turned out that the both + and - side of the output stages for CH2 were oscillating at ~600kHz. If I use a capacitance sticks to touch arbitrarily around the components, it stops their oscillation and they stay calm.
It means that the phase margin becomes small while the circuit starts up.
I decided to increase the capacitances C6 and C20 (WIMA 150pF) to 330pF (WIMA FPK2 100V) and the oscillation was tamed. 220pF didn't stop them. After visually checked the signal behavior with an oscilloscope, the unit was passed to Anchal for the TF test.
The modification was also recorded in the DCC S2100619 |
Attachment 1: PXL_20211217_001735762.jpg
|
|
Attachment 2: PXL_20211217_001719345.jpg
|
|
Attachment 3: PXL_20211217_005344828.jpg
|
|
Attachment 4: PXL_20211217_010131027.PORTRAIT.jpg
|
|
Attachment 5: PXL_20211217_011423823.jpg
|
|
Attachment 6: HAMA_Driver_V4.pdf
|
|
16515
|
Thu Dec 16 15:54:08 2021 |
Koji | Update | Electronics | ITMX feedthroughs and in-vac cables installed |
Thanks for the installation.
With regard to the connector convention, let's use the attached arrangement so that it will be consistent with the existing flange DSUB configuration. Not a big deal.
|
Attachment 1: PXL_20211216_235056582.jpg
|
|
16514
|
Thu Dec 16 15:32:59 2021 |
Anchal | Update | BHD | Finished Coil driver (odd serial number) units tests |
We have completed modifications and testing of the HAM Coil driver D1100687 units with serial numbers listed below. The DCC tree reflects these changes and tests (Run/Acq modes transfer functions).
SERIAL # |
TEST result |
S2100609 |
PASS |
S2100611 |
PASS |
S2100613 |
PASS |
S2100615 |
PASS |
S2100617 |
PASS |
S2100619 |
FAIL (CH2 phase) |
S2100621 |
PASS |
S2100623 |
PASS |
S2100625 |
PASS |
S2100627 |
PASS |
S2100629 |
PASS |
S2100631 |
PASS |
S2100633 |
Waiting for more components |
S2101649** |
PASS |
S2101651** |
PASS |
S2101653** |
PASS |
S2101655** |
PASS |
** A fix had to be done on the DC power supply for these. The units' regulated power boards were not connected to the raw DC power, so the cabling had to be modified accordingly.
Further, Paco fixed the two even serial number units (S2101648, S211650) that failed the test. The issues were minor soldering mistakes that were easily resolved. |
16513
|
Thu Dec 16 15:04:12 2021 |
Chub | Update | Electronics | ITMX feedthroughs and in-vac cables installed |
The ITMX 10" flange with four DSUB-25 feedthroughs has been install with the cables connected at the in-vac side. See photo; as requested, LO1-1 and LO1-2 are connected to the top row of feedthroughs from left to right respectively and the opposite ends of the cables placed left to right on the laser table. PR2-1 and PR2-2 are connected to the lower row of feedthroughs from left to right respectively, with the opposite ends placed on the surface below the laser from left to right. This seemed the easiest way to keep the cable orientation clear. |
Attachment 1: ITMX_feedthrough_install_12-16-21.jpg
|
|
16512
|
Thu Dec 16 12:21:16 2021 |
Anchal | Update | BHD | Coil driver test failed for S2100619-v1 |
Today I found one of the coil driver boards, S2100619 failed the test on CH2. There appears to be an extra phase lag after 10 kHz and some resonant-like feature at 7 kHz. This of course is very high-frequency stuff and maybe we don't care about these deviations. But it could mean something is off with the channel and could potentially lead to failure in the relevant frequency band in the future. I'll need help to debug this. Please see the attachment for details of test failure. |
Attachment 1: D1100687_S2100619-v1_TF_CH2_Not_Matching.pdf
|
|
16511
|
Wed Dec 15 22:09:53 2021 |
Yehonathan | Update | BHD | SOS assembly |
{Paco, Yehonathan, Anchal}
I cut the wires from the winches and removed the winch adapter plate. I engrave 'LO1' on the SOS tower. Me and Paco wrapped the SOS with foil and transported it to Anchal who put it inside the ITMX chamber.
The transportation seems to be successful. Nothing broke. However, we found that even with the short side OSEM the LO1, as it is now, cannot sit in its designed position since ITMX's side OSEM stands in its way.
If there are magnets on both ITMX sides we can move its side OSEM to the other side. Another option is to resuspend LO1 with a side magnet on its left side.
|
16510
|
Wed Dec 15 17:44:18 2021 |
Koji | Summary | BHD | Part VIII of BHR upgrade - Placed LO1 |
If ITMX already has another side magnet, we can migrate the side OSEM of ITMX to the other side. This way, the interference of the OSEMs can be avoided. |
16509
|
Wed Dec 15 16:11:38 2021 |
Anchal | Summary | BHD | Part VIII of BHR upgrade - Placed LO1 |
[Anchal, Yehonathan, Paco]
Today we opened ITMX chamber and removed the following optics and placed them in the Xend flow bench (see attachment 1):
Yehonathan brought his first SOS baby next to ITMX chamber. The suspension was carried by hands throughout. He gave me the suspension over the IMC beam tube from where I placed it on the table. I checked through the OSEMs and the face magnets were still on. I could not verify the side magnet but nothing seemed out of place.
I then moved LO1 near its planned place. I had to bolt it at 1 inch North and 0.5 inch West of its planned position because the side OSEM on ITMX is long and protrudes out of the base footprint. Even if it was small, the current layout would make the OSEM pins of the side OSEMs of ITMX and LO1 very near each other. So we can not place LO1 closer to ITMX from current position. This means the layout needs to be redesigned a bit for the modified position of LO1. I believe it will significantly shift and turn the beam from LO1 to LO2, so we might need to change the beam upstream from TT2 onwards. More discussion is required.
Unfortunately, what I thought was clicking photos was just changing modes between video and image mode, so I have no photos from today but only a video that I recorded in the end.
Photos: https://photos.app.goo.gl/23kpCknP3vz7YVrS
|
Attachment 1: signal-2021-12-15-161437.jpeg
|
|
16508
|
Wed Dec 15 15:06:08 2021 |
Jordan | Update | VAC | Vacuum Feedthru Install |
Jordan, Chub
We installed the 4x DB25 feedthru flange on the North-West port of ITMX chamber this afternoon. It is ready to go. |
16507
|
Wed Dec 15 13:57:59 2021 |
Paco | Update | Computers | upgraded ubuntu on zita |
[Paco]
Upgraded zita's ubuntu and restarted the striptool script. |
16506
|
Tue Dec 14 19:29:42 2021 |
Paco | Update | BHD | 1Y0 rack work for LO1 |
[Paco]
Two coil drivers have been installed on 1Y0 (slots 6, 7, for LO1 SOS). All connections have been made from the DAC, AI board, DAC adapter, Coil driver, Sat Amp box. Then no SOS load installed, all return connections have been made from Sat Amp box, ADC adapter, AA board, and to ADC. We will continue this work tomorrow, and try to test everything before closing the loop for LO1 suspension. |
16505
|
Tue Dec 14 14:02:33 2021 |
Yehonathan | Update | BHD | SOS assembly |
{Yehonathan, Paco}
I fixed the overall height of the adapter (attachment 1). I put an OSEM next to the side magnet. I positioned a camera in front of the SOS and connected it to my laptop for live streaming. I painted a line indicating the height of the OSEM plates and a line in between with the mean height. I discarded the wire clamp on the suspension block I released the wires from the winches and pulled on them until the magnet was roughly in the right position. I clamped the wires back on the winches and adjusted them until the magnet was on the middle line. I also verified that the roll of the adapter is aligned as before by making sure that the horizontal features on the adapter are parallel to the horizontal features on the SOS tower.
The wires were clamped to the suspension block using a new wire clamp.
I Found that locking the counterweight setscrew changes the alignment. Today we verified this effect. We released the setscrew and pre-compensated by adjusting the counterweight such that when the setscrew was locked the mirror was aligned.
We measured mechanical resonances (attachment 2). This time the yaw motion was very quiet so we got a smaller peak for the yaw. The peaks are the same as before. Y readout has peaks around the pitch and pos resonances that don't appear in the X readout. I'm not sure what they are. Maybe coming from the QPD electronics.
We locked the adapter using the EQ stops. We made sure the alignment stays close to ideal.
We installed OSEMs on the SOS. We chose suboptimal OSEMs because LO1 will only be used for steering. I made a spreadsheet copying the OSEM catalog into it. There we mark which OSEM goes where.
I cleaned the optic using the ion gun with a pressure of 30 PSI.
The next steps are:
1. Engrave the SOS tower.
2. Cut the wire at the winches and remove the winch adapter plate.
3. Wrap the SOS with foil.
4. Install the SOS in the vacuum chamber.
|
Attachment 1: magnetcentring.png
|
|
Attachment 2: FreeSwingingSpectra.pdf
|
|
16504
|
Tue Dec 14 11:33:29 2021 |
Tega | Update | PEM | git repo for temp sensor and sus medm |
[Temperature sensor]
Added new temp EPICs channels to database file (/cvs/cds/caltech/target/c1pem1/tempsensor/C1PEMaux.db)
Added new temp EPICs channels to slow channels ini file (/opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini)
[SUS medm screen]
Moved new SUS screen to location : /opt/rtcds/userapps/trunk/sus/c1/medm/templates/NEW_SUS_SCREENS
Place button on the new screen to link to the old screen and replace old screens link on sitemap.
Fixed Load Coefficient button location issue
Fixed LOCKIN flow diagram issue
Fixed watchdog labelling issue
Linked STATE WORD block to FrontEnd STATUS screen
Replaced the 2x1 pit/yaw filter screens for LOCK and DAMP fliters with 3x1 LPY filter screen
*Need some more time to figure out the OPTLEV red indicator |
16503
|
Mon Dec 13 15:05:47 2021 |
Tega | Update | PEM | git repo for temp sensor and sus medm |
[temperature sensor]
git repo: https://git.ligo.org/40m/tempsensor.git
todo
Update the temp sensor channels to fit with cds format, ie. "C1:PEM-TEMP_EX", "C1:PEM-TEMP_EY", "C1:PEM-TEMP_BS"
- Use FLOAT32_LE data format for the database file (/cvs/cds/caltech/target/c1pem1/tempsensor/C1PEMaux.db) to create the new channels.
- Keep the old datadase code and channels so we can compare with new temp channels afterwards. Also we need a 1-month overlap b4 deleting the old channels.
[sus medm screen]
git repo: https://git.ligo.org/40m/susmedmscreen.git
todo (from talk with Koji)
- Link stateword display to open "C1CDS_FE_STATUS.adl"
- Damp filter and Lock filter buttons should open a 3x1 filter screen so that the 6 filters are opened by 2 buttons compared to the old screen that has 3 buttons connected to 2X1 filter screen
- Make the LOCKIN signla modulation flow diagramlook more like the old 40m screen since that is a better layout
- Move load coefficient button to top of sus medm screen (beside stateword)
- The rectangular red outline around the oplev display is confusing and needs to be modified for clarity
- COMM tag block should not be 3D as this suggests it is a button. Make it flat and change tag name to indicate individual watchdog control as this better reflect its functionality. Rename current watchdog switch to watchdog master is it does what the 5 COMM switches do at once.
- Macro pass need to be better documented so that when we call the sus screens from locations other than sitemap, we should know what macro variables to pass in, like DCU_ID etc.
- Edit sitemap.adl to point only to the new screens. Then create a button on the new screen that points to the old screen. This way, we can still access the old screen without clogging sitemap.
- Move the new screen location to a subfolder of where the current sus screens reside, /opt/rtcds/userapps/trunk/sus/c1/medm/templates
- Rename the overview screen (SUS_CUST_HSSS_OVERVIEW.adl) to use the SUS_SINGLE nomenclature, i.e. SUS_SINGLE_OVERVIEW.adl
- Keep an eye of the cpu usage of c1pem as we add BLRMS block for other optics.
|
16502
|
Fri Dec 10 21:35:15 2021 |
Koji | Summary | SUS | Vertex SUS DAC adapter ready |
4 units of Vertex SUS DAC adapter (https://dcc.ligo.org/LIGO-D2100035) ready.
https://dcc.ligo.org/LIGO-S2101689
https://dcc.ligo.org/LIGO-S2101690
https://dcc.ligo.org/LIGO-S2101691
https://dcc.ligo.org/LIGO-S2101692
The units are completely passive right now and has option to extend to have a dewhitening board added inside.
So the power switch does nothing.
Some of the components for the dewhitening enhancement are attached inside the units.
|
Attachment 1: PXL_20211211_053155009.jpg
|
|
Attachment 2: PXL_20211211_053209216.jpg
|
|
Attachment 3: PXL_20211211_050625141-1.jpg
|
|
16501
|
Fri Dec 10 19:22:01 2021 |
Koji | Update | VAC | Pumping down the RGA section |
The scan result was ~x10 higher than the previously reported scan on 2020/9/15 (https://nodus.ligo.caltech.edu:8081/40m/15570), which was sort of high from the reference taken on 2018/7/18.
This just could mean that the vacuum level at the RGA was x10 high.
We'll just go ahead with the vacuum repair and come back to the RGA once we return to "vacuum normal".
Meanwhile, I asked Jordan to turn off the RGA to make it cool down. I shut off RGA section and turned TP2 off. |
16500
|
Fri Dec 10 18:55:58 2021 |
Tega | Update | CDS | New SUS medm screen update |
Turns out the BLRMS monitoring channels for MC1, MC2, MC3, ITMY and SRM already exist in c1pem. So I modified the new SUS screen to display the BLRMS info for the aforementioned optics. Next step is to add the BLRMS monitor for PRM, ITMX, ETMX and ETMY. This would require extending the number of inputs for the "SUS" block in c1pem to accomodate the additional inputs from the remaining optics. |
Attachment 1: BLRMS_ITMY_screenshot.png
|
|
16499
|
Fri Dec 10 15:59:23 2021 |
Paco | Update | BHD | Finished Coil driver (even serial number) units tests |
[Paco, Anchal]
We have completed modifications and testing of the HAM Coil driver D1100687 units with serial numbers listed below. The DCC tree reflects these changes and tests (Run/Acq modes transfer functions).
SERIAL # |
TEST result |
S2100608 |
PASS |
S2100610 |
PASS |
S2100612 |
PASS |
S2100614 |
PASS |
S2100616 |
PASS |
S2100618 |
PASS |
S2100620 |
PASS |
S2100622 |
PASS |
S2100624 |
PASS |
S2100626 |
PASS |
S2100628 |
PASS |
S2100630 |
PASS |
S2100632 |
PASS |
S2101648** |
FAIL (Ch1, Ch3 run mode) |
S2101650** |
FAIL (Ch3 run mode) |
S2101652** |
PASS |
S2101654** |
PASS |
** A fix had to be done on the DC power supply for these. The units' regulated power boards were not connected to the raw DC power, so the cabling had to be modified accordingly (see Attachment #1) |
Attachment 1: dc_fail.jpg
|
|
16498
|
Fri Dec 10 13:02:47 2021 |
Ian MacMillan | Summary | Computers | Quantization Noise Calculation Summary |
I am trying to replicate the simulation done by Matt Evans in his presentation (see Attachment 1 for the slide in particular).
He defines his input as so he has two inputs one of amplitude 1 at 1 Hz and one of amplitude 10^-9 at 1/4th the sampling frequency in this case: 4096 Hz
For his filter, he uses a fourth-order notch filter. To achieve this filter I cascaded two second-order notch filters (signal.iirnotch ) both with locations at 1 Hz and quality factors of 1 and 1e6. as specified in slide 13 of his presentation.
I used the same procedure outlined here. My results are posted below in attachment 2.
Analysis of results:
As we can see from the results posted below the results don't match. there are a few problems that I noticed that may give us some idea of what went wrong.
First, there is a peak in the noise around 35 Hz. this peak is not shown at all in Matt's results and may indicate that something is inconsistent.
the second thing is that there is no peak at 4096 Hz. This is clearly shown in Matt's slides and it is shown in the input spectrum so it is strange that it does not appear in the output.
My first thought was that the 4kHz signal was being entered at about 35Hz but even when you remove the 4kHz signal from the input it is still there. The spectrum of the input shown in Attachment 3 shows no features at ~35Hz.
The Input filter, Shown in attachment 4 shows the input filter, which also has no features at ~35Hz. Seeing how the input has no features at ~35Hz and the filter has no features at ~35Hz there must be either some sort of quantization noise feature there or more likely there is some sort of sampling effect or some effect of the calculation.
To figure out what is causing this I will continue to change things in the model until I find what is controlling it.
I have included a Zip file that includes all the necessary files to recreate these plots and results. |
Attachment 1: G0900928-v1_(dragged).pdf
|
|
Attachment 2: PSD_COMP_BIQ_DF2.pdf
|
|
Attachment 3: Input_PSD.pdf
|
|
Attachment 4: Input_Filter.pdf
|
|
Attachment 5: QuantizationN.zip
|